OpenStudy (anonymous):

Douglas wrapped the pots for flowers in colored foil. He wrapped 4 lily pots, 6 daffodil pots, and 5 hydrangea pots. If Douglas only wrapped one pot in green foil, what is the probability he wrapped the pot of a lily in green foil?

OpenStudy (anonymous):

4 lily pots/(4+6+5) total pots=4/15=26.667%